# SAFE: Stable Alignment Finetuning with Entropy-Aware Predictive Control for RLHF A Novel Algorithm Combining Double Soft-Min Critics with Adaptive KL Thresholding

### v1

@@ Abstract
Proximal Policy Optimization (PPO) has been positioned by recent literature as the canonical method for the RL part of RLHF. PPO performs well empirically but has a heuristic motivation and handles the KL-divergence constraint used in LM-RLHF in an ad-hoc manner and suffers form reward oscillations, entropy collapse, value function drift, and sudden policy divergence that require frequent restarts and extensive hyperparameter tuning. In this paper, we develop a new pure on policy actor-critic RL method for the LM-RLHF setting. We present SAFE (Stable Alignment Finetuning with Entropy-aware control),a novel RLHF algorithm that combines a Double Soft-Min Critic for pessimistic value estimation with a new multi-layer stabilization framework combining entropy-gated KL regulation, and PID-controlled adaptive thresholds. Unlike standard PPO’s symmetric KL penalties, SAFE distinguishes high-entropy exploration from low-entropy mode collapse and adjusts penalties dynamically based on reward velocity. Experiments on a 3B parameter model show SAFE achieves +5.15% training-average reward than PPO (0.725 vs 0.689), negligible reward crashes, and superior KL control than ppo . Our method adds minimal computational overhead and provides an interpretable, crash-resistant RLHF framework that maintains aggressive learning speed while ensuring stable long-horizon optimization suitable for production deployment. Code is available athttps://github.com/ryyzn9/SAFE
Keywords: RLHF, PPO, KL Divergence, Double Critic, Entropy Control, PID Controller
@@ 3 Method:
At the policy level, standard symmetric KL penalties fail because they penalize healthy exploratory deviations equally with dangerous overconfident drift. To address this fundamental limitation, SAFE introduces an entropy-aware predictive controller that integrates three complementary mechanisms into a unified control architecture: (1) asymmetric penalization that selectively suppresses positive KL divergence (indicating deterministic drift toward reward model artifacts) while preserving negative deviations associated with high-entropy exploration, (2) entropy-gated scaling that amplifies corrective pressure precisely when policy entropy drops into low-stochasticity regimes characteristic of mode collapse and exploitation, and (3) PID-driven adaptive thresholding that dynamically adjusts divergence tolerance based on reward velocity rather than fixed schedules. By coupling directional divergence control with entropy dynamics and temporal reward trends, this controller detects exploitative behavior earlier and maintains stable regulation across varying training phases.
@@ 3.1 Motivation: Late-Training Entropy Collapse in RLHF
A critical instability in long-horizon RLHF is late-stage policy entropy collapse: policies initially optimize successfully but undergo abrupt divergence characterized by rapid KL growth, entropy collapse, and catastrophic reward degradation. We observed the following failure pattern in a double soft critic with standered kl experiments:
@@ 3.3 Empirical Behavior
Figure 1 compares training dynamics under asymmetric control versus standard PPO. The asymmetric controller maintains KL within bounded ranges ($|\text{KL}|<0.5$ throughout 2000 steps), prevents entropy collapse (entropy >2.0>2.0), and eliminates late-stage reward degradation observed in PPO (2 crashes with $20\%$ reward loss). However, value function stability remains suboptimal (435 spikes >0.1>0.1), motivating the entropy-aware predictive controller introduced in Section 3.4.
@@ 3.4.3 Reward-Driven Predictive Threshold Adaptation
Rather than using a fixed KL target, the controller adapts its divergence tolerance using a PID controller driven by reward feedback.
Let $\text{PID}(r_{t})$ denote the proportional–integral–derivative update computed from the reward signal. The adaptive KL threshold is defined as:
$\tau_{t}=(\tau_{\text{base}}+\text{PID}(r_{t}))\cdot\phi_{t},$ (11)
where:
  • $\tau_{\text{base}}$ is the baseline divergence tolerance,
  • $\phi_{t}$ is a multiplicative factor provided by the phase detector,
  • the PID controller adjusts sensitivity based on recent reward dynamics.
This design allows the controller to relax divergence constraints during productive optimization phases and tighten them when reward improvement stagnates or destabilizes.
@@ 3.4.7 Controller Algorithm
Algorithm 2 Entropy-Aware Predictive Controller (Plan D)
@@
3.4.8 Practical Role in SAFE
@@ 4.3.2 Entropy-Gated Penalty
4.4 PID-Controlled Adaptive Threshold
To adapt divergence tolerance during training, a PID controller dynamically adjusts the KL threshold.
@@ 4.4 PID-Controlled Adaptive Threshold
$\tau_{t}=\tau_{\text{base}}+K_{p}e_{t}+K_{i}\sum_{k=0}^{t}e_{k}+K_{d}(e_{t}-e_{t-1}),$ (25)
@@
Here:
  • $\text{EMA}(r_{t})$ denotes exponentially smoothed reward,
  • $v_{\text{target}}$ is a fixed small positive target improvement rate,
  • $K_{p},K_{i},K_{d}$ are PID gains.
@@
which prevents overly aggressive or overly permissive divergence constraints.
@@ 4.5 Phase Detection Algorithm
Training dynamics differ substantially between early exploration, reward climbing, plateau phases, and late convergence. To adapt divergence tolerance across these regimes, we implement a lightweight phase detector that monitors reward statistics over sliding windows and applies multiplicative adjustments $\phi_{t}$ to the KL threshold.
Phase Detection Algorithm. The detector maintains a 100-step reward history and classifies the current training phase based on mean reward trends and variance:
@@
The phase multipliers directly modulate the adaptive threshold $\tau_{t}=(\tau_{\text{base}}+\text{PID}(r_{t}))\cdot\phi_{t}$ as described in Equation (25). This design enables:
@@ 5 Training Algorithm
6 Objective Function
$L_{total}=L_{PPO}+0.5L_{Huber}+L_{KL}-\beta H(\pi_{\theta})$ (32)
Where:
$L_{PPO}=-\mathbb{E}[\min(\rho_{t}A_{t},clip(\rho_{t})A_{t})]$ (33)
$L_{Huber}=\frac{1}{2}[Huber(V_{soft},r)+Huber(V_{clip},r)]$ (34)
@@ Appendix A Theoretical Analysis: Hypothesized Reward Hacking Resistance
A rigorous assessment of SAFE’s reward-hacking behavior requires dedicated experiments with explicitly hackable reward functions, adversarial and out-of-distribution prompts, and long training horizons. Designing and running such evaluations is an important direction for future work and is beyond the scope of the present study.

### v2

# SAFE: Stable Alignment Finetuning with Entropy-Aware Predictive Control for RLHF Combining Double Soft-Min Critics with Adaptive KL Thresholding

@@ Abstract
Proximal Policy Optimization (PPO) has been positioned by recent literature as the canonical method for the RL part of Reinforcement Learning from Human Feedback (RLHF). PPO performs well empirically but has a heuristic motivation and handles the KL-divergence constraint used in LM-RLHF in an ad-hoc manner and suffers from reward oscillations, entropy collapse, value function drift, and sudden policy divergence that require frequent restarts and extensive hyperparameter tuning. In this paper, we develop a new pure on policy actor-critic RL method for the LM-RLHF setting. We present SAFE (Stable Alignment Finetuning with Entropy-aware control),a novel RLHF algorithm that combines a Double Soft-Min Critic for pessimistic value estimation with a new multi-layer stabilization framework combining entropy-gated KL regulation, and Proportional-Integral-Derivative (PID)-controlled adaptive thresholds. Unlike standard PPO’s symmetric KL penalties, SAFE distinguishes high-entropy exploration from low-entropy mode collapse and adjusts penalties dynamically based on reward velocity. Experiments on a 3B parameter model show SAFE achieves +5.15% training-average reward than PPO (0.725 vs 0.689), negligible reward crashes, and superior KL control than ppo . Our method adds minimal computational overhead and provides an interpretable, crash-resistant RLHF framework that maintains aggressive learning speed while ensuring stable long-horizon optimization suitable for production deployment. Code is available at https://github.com/ryyzn9/SAFE
Keywords: RLHF, PPO, KL Divergence, Double Critic, Entropy Control, Proportional-Integral-Derivative (PID)-controlled
@@ 3 Method:
At the policy level, standard symmetric KL penalties fail because they penalize healthy exploratory deviations equally with dangerous overconfident drift. To address this fundamental limitation, SAFE introduces an entropy-aware predictive controller that integrates three complementary mechanisms into a unified control architecture: (1) asymmetric penalization that selectively suppresses positive KL divergence (indicating deterministic drift toward reward model artifacts) while preserving negative deviations associated with high-entropy exploration, (2) entropy-gated scaling that amplifies corrective pressure precisely when policy entropy drops into low-stochasticity regimes characteristic of mode collapse and exploitation, and (3) Proportional-Integral-Derivative (PID)-driven adaptive thresholding that dynamically adjusts divergence tolerance based on reward velocity rather than fixed schedules. By coupling directional divergence control with entropy dynamics and temporal reward trends, this controller detects exploitative behavior earlier and maintains stable regulation across varying training phases.
@@ 3.1 Motivation: Late-Training Entropy Collapse in RLHF
A critical instability in long-horizon RLHF is late-stage policy entropy collapse: policies initially optimize successfully but undergo abrupt divergence characterized by rapid KL growth, entropy collapse, and catastrophic reward degradation. We observed the following failure pstandard KLdouble soft critic with standered kl experiments:
@@ 3.3 Empirical Behavior
Figure 1 compares training dynamics under Double Soft-Min Critics with Asymmetric KL control versus standard PPO over 2000 steps. The asymmetric controller demonstrates improved KL regulation but reveals significant trade-offs in other stability metrics.
KL Divergence Control. The Asymmetric KL controller achieves substantially better KL regulation, with mean absolute KL divergence of $0.241$ compared to PPO’s $0.131$. More importantly, only $28.6\%$ of training steps exceed $\text{KL}>0.2$ under Asymmetric KL versus $55.9\%$ for PPO, and KL volatility is reduced by $44\%$ (rolling std $0.294$ vs $0.526$). This indicates that directional divergence penalties successfully constrain large positive KL excursions associated with exploitative drift.
Entropy Preservation. The asymmetric mechanism maintains moderate entropy levels throughout training, with mean policy entropy remaining stable. However, completion length variability is reduced (std $9.7$ vs $12.6$ tokens), suggesting somewhat more deterministic generation patterns.
Reward Performance. Despite improved KL control, Double Soft-Min Q with Asymmetric KL achieves lower mean reward than PPO ($0.672$ vs $0.689$, $p<10^{-15}$), representing a $2.5\%$ performance gap. Both methods exhibit 2 reward crashes exceeding $20\%$ loss, indicating that asymmetric penalties alone do not eliminate catastrophic failures. Additionally, reward volatility is $1.6\times$ higher under Asymmetric KL control (rolling std $0.0343$ vs $0.0208$), and the reward coefficient of variation is $0.081$ compared to PPO’s $0.114$.
Value Function Instability. The most severe limitation is critic destabilization: the Double Soft-Min Q with Asymmetric KL produces 1256 value loss spikes exceeding $0.1$, compared to only 28 under PPO—a $45\times$ increase in instability. Mean value loss rises to $0.226$ versus PPO’s $0.006$, suggesting that the interaction between asymmetric penalties and advantage estimation introduces severe optimization noise. This critic instability likely explains the reduced reward performance and higher training volatility.
These results demonstrate that while Asymmetric KL control successfully constrains divergence magnitude, it destabilizes value learning and sacrifices reward performance. The directional penalty approach combined with pessimistic double critics alone is insufficient for robust RLHF optimization. These findings motivate the entropy-aware predictive controller introduced in Section 3.4, which addresses value instability through dual-timescale tracking, adaptive thresholding, and phase-aware modulation to coordinate policy and critic stabilization.
@@ 3.4.3 Reward-Driven Predictive Threshold Adaptation
Rather than using a fixed KL target, the controller adapts its divergence tolerance using a PID controller driven by reward velocity feedback. This enables the system to automatically adjust exploration constraints based on whether training is improving at the desired rate.
The PID controller tracks the reward improvement rate rather than absolute reward magnitude. The control error is defined as:
$e_{t}=\Delta\text{EMA}(r_{t})-v_{\text{target}},$ (11)
where $\Delta\text{EMA}(r_{t})=\text{EMA}_{t}-\text{EMA}_{t-1}$ is the velocity (rate of change) of the exponentially smoothed reward with decay factor $0.95$, and $v_{\text{target}}=0.001$ is the target improvement rate (0.1% per step).
The adaptive KL threshold is computed using standard PID control:
$\tau_{t}=\left(\tau_{\text{base}}+K_{p}e_{t}+K_{i}\sum_{k=0}^{t}e_{k}+K_{d}(e_{t}-e_{t-1})\right)\cdot\phi_{t},$ (12)
where $\tau_{\text{base}}$ is the baseline divergence tolerance, $K_{p},K_{i},K_{d}$ are proportional, integral, and derivative gains, and $\phi_{t}$ is the phase-dependent multiplier. The integral term is clamped to [1,1][-1,1] to prevent windup. The resulting threshold is bounded to $[0.1,0.6]$ to maintain reasonable operating ranges.
The mechanism works as follows: When reward improvement exceeds the target rate ($e_{t}>0$), the PID output becomes positive, increasing $\tau_{t}$ and allowing greater KL divergence to encourage continued exploration of the productive region. When improvement falls below target ($e_{t}<0$)—whether due to stagnation or decline—the PID output becomes negative, decreasing $\tau_{t}$ and tightening KL constraints to prevent drift and stabilize learning. The proportional term responds to the current velocity gap, the integral term accumulates persistent under-performance to force corrective action during prolonged stagnation, and the derivative term responds to acceleration changes to preemptively adjust before large deviations occur. The phase multiplier $\phi_{t}$ further modulates this behavior across training regimes (warmup, climbing, plateau, converged).
This design creates a feedback loop where divergence tolerance tracks optimization productivity rather than absolute performance, relaxing constraints during sustained improvement and tightening them during plateaus or instability.
@@ 3.4.7 Controller Algorithm
Algorithm 2 Entropy-Aware Predictive Controller
@@
3.4.8 Practical Role of the Entropy-Aware Controller in SAFE
@@ 4.3.2 Entropy-Gated Penalty
4.4 Proportional-Integral-Derivative (PID)-controlled Adaptive Threshold
To adapt divergence tolerance dynamically during training, we employ a PID controller that adjusts the KL threshold based on training progress. Unlike fixed schedules that cannot respond to varying learning dynamics, the PID controller tracks reward velocity and modulates exploration constraints accordingly.
@@ 4.4 Proportional-Integral-Derivative (PID)-controlled Adaptive Threshold
$\tau_{t}=\left(\tau_{\text{base}}+K_{p}e_{t}+K_{i}\sum_{k=0}^{t}e_{k}+K_{d}(e_{t}-e_{t-1})\right)\cdot\phi_{t},$ (26)
@@
Here, $\text{EMA}(r_{t})$ denotes exponentially smoothed reward with decay factor $0.95$, $\Delta\text{EMA}(r_{t})=\text{EMA}_{t}-\text{EMA}_{t-1}$ is the velocity (improvement rate), $v_{\text{target}}=0.001$ is the target improvement rate, $K_{p},K_{i},K_{d}$ are proportional, integral, and derivative gains, and $\phi_{t}$ is the phase-dependent multiplier (Section 4.5).
The three PID components serve complementary roles:
  • Proportional term ($K_{p}e_{t}$): Responds immediately to the current velocity gap, increasing $\tau_{t}$ when improvement exceeds target and decreasing it during stagnation.
  • Integral term ($K_{i}\sum e_{k}$): Accumulates persistent under-performance, forcing corrective tightening during prolonged plateaus. The sum is clamped to [1,1][-1,1] to prevent integral windup.
  • Derivative term ($K_{d}(e_{t}-e_{t-1})$): Responds to acceleration changes, preemptively adjusting before large deviations occur.
@@
which prevents overly aggressive constraints (that would halt exploration) or overly permissive bounds (that would allow uncontrolled drift). This creates a feedback loop where divergence tolerance tracks optimization productivity: relaxing constraints during sustained improvement and tightening them during plateaus or instability.
@@ 4.5 Phase Detection Algorithm
Training dynamics differ substantially between early exploration, reward climbing, plateau phases, and late convergence. To adapt divergence tolerance across these regimes, we implement a lightweight phase detector that monitors reward statistics over sliding windows and applies multiplicative adjustments $\phi_{t}$ to the KL threshold. Phase Detection Algorithm. The detector maintains a 100-step reward history and classifies the current training phase based on mean reward trends and variance:
@@
The phase multipliers directly modulate the adaptive threshold $\tau_{t}=(\tau_{\text{base}}+\text{PID}(r_{t}))\cdot\phi_{t}$ as described in Equation (4.4). This design enables:
@@ 5 Training Algorithm
6 Overall Objective Function
SAFE optimizes a composite loss balancing policy improvement, value learning, divergence control, and exploration:
$L_{\text{total}}=L_{\text{PPO}}+0.5L_{\text{Huber}}+L_{\text{KL}}-\beta H(\pi_{\theta}).$ (33)
PPO Policy Loss.
The clipped surrogate objective Schulman et al. (2017):
$L_{\text{PPO}}=-\mathbb{E}\left[\min\!\left(\rho_{t}\hat{A}_{t},\,\text{clip}(\rho_{t},1-\epsilon,1+\epsilon)\hat{A}_{t}\right)\right],$ (34)
with importance ratio $\rho_{t}=\pi_{\theta}(a_{t}\mid s_{t})/\pi_{\text{old}}(a_{t}\mid s_{t})$ and advantage $\hat{A}_{t}$ from the pessimistic soft-min critic.
Value Loss.
Clipped Huber regression Huber (1964) for robust critic learning:
[…]
which caps gradient magnitude from reward-model outliers.
KL Penalty.
Entropy-gated, adaptively thresholded divergence control (Section 4.3):
$L_{\text{KL}}=\begin{cases}0,&\bar{D}_{t}\leq\tau_{t},\\ \lambda(\bar{D}_{t}-\tau_{t})^{2}\cdot\max\!\left(0.5,\frac{H_{\text{floor}}}{H(\pi_{\theta})+\epsilon}\right),&\text{otherwise},\end{cases}$ (36)
with PID-adjusted threshold $\tau_{t}$ and entropy-dependent scaling.
Entropy Bonus.
Standard term $\beta H(\pi_{\theta})$ maintains baseline exploration.
This objective extends PPO with pessimistic value aggregation, Huber-based robustness, and adaptive entropy-aware KL control for stable long-horizon RLHF.
@@ Appendix A Theoretical Analysis: Hypothesized Reward Hacking Resistance
A rigorous assessment of SAFE’s reward-hacking behavior requires dedicated experiments with explicitly hackable reward functions, adversarial and out-of-distribution prompts, and long training horizons. Designing and running such evaluations are an important direction for future work and is beyond the scope of the present study.